# AI Competitions and Benchmarks: towards impactful challenges with post-challenge papers, benchmarks and other dissemination actions

### v2

@@ Abstract
Organising an AI challenge does not end with the final event. The long-lasting impact also needs to be organised. This chapter covers the various activities after the challenge is formally finished. The target audience of different post-challenge activities is identified. The various outputs of the challenge are listed with the means to collect them. The main part of the chapter is a template for a typical post-challenge paper, including possible graphs as well as advice on how to turn the challenge into a long-lasting benchmark.
This document is the 5th chapter of the book : AI Competitions and Benchmarks, the science behind the contests https://sites.google.com/chalearn.org/book/home
@@
If winners are announced at the end of a competition, and everyone returns to their usual activity, even the best thought competition would have been useless. This chapter covers the various post-challenge activities necessary for the long-lasting impact of a challenge, some of which should be prepared already before the challenge.
@@
Four types of stakeholders play specific roles in a challenge: the organisers, the application domain experts, the AI/Machine Learning experts on techniques relevant to the problem posed, and the participants. The participants are diverse; they can be experts from the application domain or AI, students or seasoned data scientists. Input from all these people should be gathered and made available for posterity.
This chapter is organised as follows: Sec. 4.1 is an introduction, Sec. 4.2 details the raw output expected from a challenge, while Sec. 4.3 deals with post-challenge workshops and discussions. Sec 4.4 proposes a typical post-challenge template, and Sec 4.5 indicates how to build long-lasting benchmarks before the conclusion Sec. 4.6.
4.1 Purpose of post-challenge activity
The main purpose of a challenge is to address an open problem and engage a whole community towards solving it, not only at some point in time but hopefully in the long run. To keep such a community engaged and hopefully growing, it is important to share with them and the wider research community:
  • the results and lessons on which they can build upon,
  • the remaining gaps/challenges and related most essential research directions to aim at,
  • any materials that could facilitate further research, such as benchmarks, solution codes, visualisation and analysis posts or projects, tutorial posts or videos.
@@ 4.1 Purpose of post-challenge activity
To this end, post-challenge activity is necessary, especially as challenges raw and meta outputs are often numerous. This activity should be structured to assess the potential value of those different outputs and eventually best organise them to extract meaningful analysis, discussion and conclusions. Sufficient resources must be assigned upfront to this activity.
A post-challenge paper, which we will discuss, is a valuable milestone to aim at to conduct this activity in a structured fashion, clarifying the results, lessons, gaps and research directions in a concise and intelligible form. It further helps keep the community engaged by disseminating challenge outcomes and giving incentives to advance state of the art further. Note that there should ideally exist a white paper addressing the public at large and delivering the necessary background and concepts to understand the problem, a challenge design paper describing the modelling and problem implementation, and any code or dataset documentation.
As for post-challenge papers, they are often aimed at a more technical audience. Both challenge organisers and participants can contribute to such papers. In the end, there can be many different paper kinds such as:
@@
  • A short analysis paper by organisers only (based fact-sheet) comparing best solution performance and reflecting on challenge design
@@
  • A federated paper including best participants with a stronger focus on best solution descriptions
@@
  • An introduction to a book or special issue for several participant papers
  • An introduction to proceedings of a workshop
@@
  • A journal paper possibly following several iterations of a competition series with a more in-depth analysis of how the challenge eventually altered the scientific landscape and an emphasis on the current applicable state of the art.
A post-challenge paper will also offer variations depending on the related challenge objectives: was it addressing a fundamental Machine Learning scientific question or an applied problem? Was it a new problem or a new formulation of an existing one? Was it mostly about assessing the problem feasibility currently or evaluating and advancing more at large existing state of the art?
@@
When starting to write such a paper, one should consider the diversity of possible readers and decide how to structure and position it best. Different readerships can be anticipated:
@@
  • AI/ML researchers looking to apply their approaches to various problems (Federated paper or introduction to proceedings of a workshop )
@@
  • AI/ML researchers looking to know more about the state-of-the-art methods (federated paper, introduction to special issue/book or journal paper)
@@
  • Domain Expert Researcher interested in the outcomes relevant to their community and possible comparison with non-AI methods (short or journal paper)
  • Domain Expert Scientist looking for problem modelling and how to best frame them to engage the AI community (short paper)
  • Other challenge organisers looking for any best practices and innovations on how to design, set up and run successfully such challenges (short or journal paper)
  • Scientist looking for evaluation framework definition and design (short or journal paper)
@@
  • Research managers looking for fruitful collaboration with skilful teams (federated paper)
@@
  • Vendors/Investors looking for promising approaches in nextgen applications (short or journal paper)
  • Science popularisers/reporters (short or journal paper
inline,color=brown!40]It would be useful to reference to papers of the nature referred to in the itemisation, and comment on why such paper could be prototypical. Otherwise this is very dry and theoretical enumeration of classes of papers… – Gustavo inline,color=cyan!40]In principle yes… – David
@@
Having those high-level considerations in mind and before describing typical post-challenge papers more concretely, we will now review the different kinds of challenge outputs that could be made available and of good use.
4.2 Challenge raw output
The challenge raw output is all the material produced during the challenge, which must be analysed. The various types of outputs are listed in this section.
4.2.1 Competition platform output
A typical challenge platform can provide to the organisers, for each participant (or team) and for all their submissions, many pieces of information: the time of submission, public and private scores and other similar quantities and most likely the detailed content of the submission, including the actual code in case of code submissions. The final (or selected) submissions are the most interesting, but intermediate ones can also inform the improvement process. Also, the time evolution of the public and private leaderboards is part of the narrative of the challenge.
4.2.2 Participants fact sheet
At the end of the competition, it is good practice for organisers to submit a form, the “fact sheet”, to the participants to (i) have details on their best submission, which is not automatically provided by the platform, and (ii) know more about the participant themselves, who they are and how they managed their participation. It is best to advertise the form already in the last few days of the competition before participants move on to other activities and to insist that all inputs are worthwhile, even from non-top performers. The form should have a good balance between closed form Multiple Choice Questions, more straightforward to analyse, and free fields to gather specific feedback. The form should be anonymous by default but with the possibility of indicating the platform user name or leaving an email to continue the dialogue. The goal is to provide an overview of all participants (at least the ones who answer), while the bulk of the post-challenge activity will (and rightly so) focus on the best or most original contributions.
@@ 4.2.2 Participants fact sheet
  • Which available material and resources (papers, documentation, tutorials, baselines, tools, compute credits) did participants know of, and how useful and easy to use were they? This could help understand if the main competition features were understood or if any were missed if that helps participants stay active and engaged, and if everything was eventually there to help them learn and participate. This would help organisers improve those resources for future benchmark or competition iterations.
@@
4.2.3 Code
In a code-less platform, participants train their algorithm on a training dataset, apply it (the inference part) to a test dataset and submit the results as a data file. In this case, asking participants to submit their software, including training and inference, is customary on a platform like GitHub. This, however, can only be made mandatory for participants who can claim a price, which should have been specified in the competition rules; the others would often not bother. Also, if the inference code can be tested relatively easily, it is much less the case for the training code, as the final model is often the result of many iterations. The code should be runnable, well documented, and accompanied by a short document describing its functionality.
@@ 4.2.4 Competition log
4.2.5 Unorganized raw output
A flow of information from the competition is not formatted by the platform or online forms and needs to be analysed. They can also represent a measure of how active the competition is. The competition forum is the primary source of such information: data exploration posts and notebooks, insights, code sharing, and documentation gleaned on the web on the topic. One can, for example, see an idea appear in a forum post, followed by a code implementation (not necessarily by the same person), followed by a general increase in the scores on the leaderboard. However, relevant technical discussions may happen outside the competition forum: on social networks, blog posts or arXiv papers. Also, competitions are often used as practical projects for (e.g., data science) courses. Compared to typical participants, students are often compelled to write a hopefully clear document about their techniques, which might be public. Such spontaneous output should be harvested regularly, which is much easier if the competition has a unique acronym to be googled and for which alerts can be set up.
4.3 Post-challenge workshops and discussion
One or more post-challenge workshops are often organised as part of a larger conference or in a dedicated venue. Participating in it is part of the incentive for participants. It allows participants to switch from competition to cooperation. They can discuss their techniques between them and with the application or Machine Learning experts. The connections created during such workshops encourage participants to remain engaged with the competition topic instead of just moving on to a different one. The discussions during the workshops and continued online are key to deriving the scientific lessons from the workshops and the avenues for further study and possible future challenges.
@@ 4.3 Post-challenge workshops and discussion
4.4 Post challenge paper template
@@ 4.4.1 Introduction
4.4.2 Challenge Description
This section should ideally come as a reminder and synthesis of a prior paper on the challenge design. One should first review the problem of the challenge to give a good enough understanding to the targeted audience for the remainder of the paper, particularly highlighting its main dimensions of complexity and variations. One should further present the specific task formulated for the challenge with some description of the underlying datasets, framework (such as a chatbot or reinforcement learning framework if relevant) and problem modelisation (possibly relying on some simulator). One can remind if the challenge aims at delivering a new problem or a new formulation to the Machine Learning community or advancing and benchmarking state-of-the-art. A section on related works and similar challenges can be written to best position this challenge in the scientific landscape and possibly build upon previous work.
The scoring metric and protocol should be discussed, sharing some considerations as to why they were chosen a priori to evaluate any advances towards problem-solving best. A brief description of the challenge platform , as well as possible specific choices (like resource allocation) or developments that can be justified there.
@@ 4.4.3 Challenge Narrative
For example, Fig. 4.1 shows the progress of participants in one competition as a function of time; one can see the bulk of participants progressing as a swarm, following community understanding of the problems, while a few isolated outliers obtained the best scores.
Figure 4.1: Participants’ best score evolution as a function of days in the competition TrackMLAccuracy2019 .
@@
Fig. 4.2 shows participants’ progress in a different competition in the (accuracy, speed) plan. Various strategies are evident: some tried to optimise both simultaneously, why some others (the best), fastrack and gorbuno have first reached the largest accuracy, then improved their speed without accuracy loss.
@@
Figure 4.2: Participants score evolution: the horizontal axis is the accuracy, and the vertical axis is the inference speed. The total score, a function of both variables, is displayed in grey contours. Each colour/marker type corresponds to a contributor; the lines help to follow the score evolutionTrackMLThroughput2021 .
@@
A graph like Fig. 4.3 can help summarise the main events and competition activity.
@@
Figure 4.3: A high-level retrospective competition period timeline can help understand the competition organisation through phases and noteworthy events such as competition adjustments, peak of activities, performance outbreaks, and collaborative periods. This can support the competition narrative. marot2021learning
4.4.4 Post challenge checks
@@
In most cases, the performance of submissions can be evaluated against a held-out (or private) dataset, providing the final ranking. Generally, this is automatically done on the challenge platform, but it may require some manual actions, particularly when doing it automatically requires too many resources. The stability of the ranking can then be evaluated as in 4.4. The graph shows that the rankings among teams are stable, except for 4th Rek, who performs significantly worse in the second phase.
@@
Figure 4.4: Overfitting/Ranking stability plot from autoseries . The comparison of the x/y axes shows overfitting. In a challenge of two phases, the feedback (or public) phase is the first phase, and the private phase is the second. By showing how submissions perform in the consequent two phases, we demonstrate the overfitting of algorithms. The rectangle shows ranking stability around each team. This rectangle is calculated by the average ranks of multiple reproductions of submissions’ performances.
@@
A more detailed analysis is also possible as in 4.5. The comparison of the private curves shows that Gabor (the winner) is clearly above the others, as confirmed by a numerical analysis. The comparison of private and public curves of different participants shows a clear overfitting case for Lubozs, who was first on the final public leaderboard but slipped to seventh position in the private leaderboard. It turned out that he had indicated in his blog that he had set up an automatic cron job, which was automatically re-submitting new submissions (five times per day, which was the maximum allowed by the platform) with slightly altered parameters to maximise his public score. This engineering feat allowed him to select a lucky spike and grab the top of the public leaderboard, but this did not fool the private leaderboard measurement. It should be noted that this piece of information was not reported through the standard channels or forum; it was discovered by googling, illustrating the point of analysing unorganised information as discussed in Sec. 4.2.5.
@@
Figure 4.5: AMS Significance, which is a figure of merit of a classifier in a physics particle discovery context, as a function of decision threshold for different submissions in the HiggsML challenge higgsml . The submission score is the maximum of the curve. (d) shows the private curves for different participants. (a),(b) and (c) compare the public and private curves from three participants, Gabor, Choko and Lubozs, with dots indicating their maximum values. The private curves are smoother because they are evaluated on more examples.
@@
Additional graphs can be produced from by-products of solution submission gathered by the platform. For example, Fig.4.6 compares the performance on different datasets, giving for each the intrinsic and modelling difficulties. Ideally, organisers should choose datasets of low intrinsic difficulty and high modelling difficulty, which would be in this example Isaac2, Caucase or freddy.
@@
Figure 4.6: Measurement of task difficulty from autocv . Each row corresponds to one dataset. Two difficulties are calculated here: intrinsic difficulty (maximum score minus best participant’s score), shown with the green bar and modelling difficulty (best participant’s score minus baseline score), shown with the orange bar.
@@
One can also analyse the solutions to evaluate their originality. For example, Fig 4.7 shows the dendrogram of a clustering competition. The dendrogram shows how similar or not the clusters found are, which correlates well with what is known on the participant’s method or background: the set of six participants at the top of the diagram (#$\#12$ to #$\#9$) have highly optimised the starting kit using generic clustering algorithms; #$\#3$ and #$\#4$ are domain experts with the same background, who have found similar clusters with optimised domain methods; #$\#1$ is a CS student who has spent a lot of time studying the domain literature, has seen similar clusters as #$\#3$ and #$\#4$; on the other hand, the diagram sets apart #$\#2$ who is a CS expert who has developed a very original approach (which turned out to be impractical because of the large resource it requires).
[…]
4.4.5 Deeper analysis of the submissions

### v5

# Towards impactful challenges: post-challenge paper, benchmarks and other dissemination actions

@@ Abstract
The conclusion of an AI challenge is not the end of its lifecycle; ensuring a long-lasting impact requires meticulous post-challenge activities. The long-lasting impact also needs to be organised. This chapter covers the various activities after the challenge is formally finished. This work identifies target audiences for post-challenge initiatives and outlines methods for collecting and organizing challenge outputs. The multiple outputs of the challenge are listed, along with the means to collect them. The central part of the chapter is a template for a typical post-challenge paper, including possible graphs and advice on how to turn the challenge into a long-lasting benchmark.
Keywords: post-challenge, analysis, paper, dissemination
1 Introduction
@@
If winners are announced at the end of a competition, and everyone simply returns to their usual activity, even the best-conceived competition would have been of limited use. Indeed, many AI/ML challenges shine briefly and are then forgotten, leading to missed opportunities for long-term impact. This chapter discusses the various post-challenge activities necessary to ensure lasting effects from such events, some of which should be prepared even before the challenge starts.
AI/ML challenges have become increasingly popular over the last decade, engaging not only researchers but also enthusiasts and industry practitioners. According to the public data from platforms such as MLContests, [^1] https://mlcontests.com/ which gathers information of challenges across platforms, thousands of competitions solving real-world or academia problems are hosted each year. Conferences like NeurIPS, and KDD have also introduced official competition tracks in recent years [^2] https://neurips.cc/Conferences/2024/CompetitionTrack [^3] https://kdd.org/kdd-cup, aiming to foster collaboration between experts and provide tangible benchmarks for new methodologies. Blog posts from conference organizers have explained their rationale for hosting more challenges, typically highlighting their value in addressing open problems, engaging communities, and producing state-of-the-art solutions [^4] https://blog.neurips.cc/2024/06/04/neurips-2024-competitions-announced/.
@@
Four types of stakeholders typically play specific roles in a challenge: the organisers, the application domain experts, the AI/Machine Learning experts on techniques relevant to the problem posed, and the participants. Participants themselves come from diverse backgrounds—ranging from domain-specific experts and AI specialists to students and experienced data scientists—each contributing unique perspectives that enrich the challenge outcomes. Input from all these people should be gathered and made available for posterity so that the community can build upon results and lessons, identify remaining gaps/research directions, and access new materials (benchmarks, codes, tutorials) to further the state of the art.
1.1 Why Post-Challenge Activities Matter
Ensuring long-term community engagement requires sharing outcomes and resources with both participants and the wider research community:
  • Results and lessons on which they can build upon,
  • Remaining gaps and research directions that can push the boundaries of current knowledge,
  • Materials (benchmarks, solution codes, visualizations, tutorials) that facilitate continued and reproducible research.
@@ 1.1 Why Post-Challenge Activities Matter
Post-challenge activity is therefore necessary, especially as the raw and meta outputs of challenges can be numerous and complex. Assigning sufficient resources for structuring and evaluating these outputs helps extract meaningful analysis, discussion, and conclusions.
A post-challenge paper represents one cornerstone of such activities. It conveys results, lessons, gaps, and research directions in a concise, intelligible form. Such a paper often keeps the community engaged by disseminating challenge outcomes, contributing new insights, and incentivizing future work to push the state of the art. Ideally, this post-challenge paper is complemented by:
  • A white paper for a broader audience, explaining the challenge background,
  • A challenge design paper describing the modelling and problem implementation,
  • Code or dataset documentation for reproducibility and further experimentation.
1.2 Types of Post-Challenge Papers
In practice, many different types of post-challenge papers are possible, each providing distinctive benefits:
@@ 1.2 Types of Post-Challenge Papers
  • A short analysis paper by organisers only (based on a fact sheet) comparing the best solutions’ performance and reflecting on challenge design.
@@
  • A federated paper that includes top participants, with a stronger focus on best solution descriptions.
@@
  • An introduction to a book or special issue that collates multiple participant papers.
  • An introduction to the proceedings of a workshop.
@@
  • A journal paper, possibly following several iterations of a competition series, providing a more in-depth analysis of how the challenge altered the scientific landscape and emphasizing the applicable current state of the art.
The exact nature of a post-challenge paper depends on the competition’s objectives: addressing a fundamental AI/ML question, or tackling an applied problem, or introducing a novel formulation of an existing issue, or focusing on feasibility, benchmarking, and pushing the current state of the art.
@@
When planning such a paper, organizers should consider the diversity of possible readers and carefully position the content. Different audiences may include:
@@
  • AI/ML researchers looking to apply their approaches to various problems (federated paper or introduction to workshop proceedings),
@@
  • AI/ML researchers wanting to learn about state-of-the-art methods (federated paper, introduction to special issue/book, or journal paper),
@@
  • Domain expert researchers interested in outcomes relevant to their field and comparisons with non-AI methods (short or journal paper),
  • Domain expert scientists seeking models and problem framings that best attract the AI community (short paper),
  • Challenge organizers searching for best practices and innovations in setting up and running competitions (short or journal paper),
  • Scientists investigating evaluation frameworks (short or journal paper),
@@
  • Research managers looking for fruitful collaboration with skilled teams (federated paper),
@@
  • Vendors/investors seeking promising approaches in next-generation applications (short or journal paper),
  • Science popularizers/reporters (short or journal paper).
@@
The chapter guides readers through critical aspects of post-challenge planning: organizing raw outputs (Sec. 2), facilitating workshops (Sec. 3), drafting post-challenge papers (Sec 4), and establishing enduring benchmarks (Sec 5), concluding with actionable recommendations in (Sec. 6).
2 Challenge raw output
Having the above mentioned high-level considerations in mind and before describing typical post-challenge papers more concretely, we will now review the different kinds of challenge outputs that could be made available and of good use. The raw output of the challenge is all the material produced during the challenge, which must be analysed. The various types of outputs are listed in this section.
2.1 Competition platform output
A typical challenge platform can provide to the organisers, for each participant (or team) and all their submissions, many pieces of information: the time of submission, public and private scores and other similar quantities and most likely the detailed content of the submission, including the actual code in case of code submissions. The final (or selected) submissions are the most interesting, but intermediate ones can also inform the improvement process. Also, the time evolution of the public and private leaderboards is part of the challenge narrative.
2.2 Participants fact sheet
At the end of the competition, it is good practice for organisers to submit a form, the “fact sheet”, to the participants to (i) have details on their best submission, which is not automatically provided by the platform, and (ii) know more about the participant themselves, who they are and how they managed their participation. It is best to advertise the form already in the last few days of the competition before participants move on to other activities and to insist that all inputs are worthwhile, even from non-top performers. The form should have a good balance between closed form Multiple Choice Questions, more straightforward to analyse, and free fields to gather specific feedback. The form should be anonymous by default, but it should also include the possibility of indicating the platform user name or leaving an email to continue the dialogue. The goal is to provide an overview of all participants (at least the ones who answer), while the bulk of the post-challenge activity will (and rightly so) focus on the best or most original contributions.
@@ 2.2 Participants fact sheet
  • Which available material and resources (papers, documentation, tutorials, baselines, tools, compute credits) did participants know of, and how useful and easy to use were they? This could help understand if the main competition features were understood or if any were missed, if that helps participants stay active and engaged, and if everything was eventually there to help them learn and participate. This would help organisers improve those resources for future benchmark or competition iterations.
@@
2.3 Code
In a code-less platform, participants train their algorithm on a training dataset, apply it (the inference part) to a test dataset and submit the results as a data file. In this case, asking participants to submit their software, including training and inference, is customary on a platform like GitHub. This, however, can only be made mandatory for participants who can claim a price, which should have been specified in the competition rules; the others would often not bother. Also, if the inference code can be tested relatively quickly, it is much less the case for the training code, as the final model is often the result of many iterations. The code should be runnable, well documented, and accompanied by a short document describing its functionality.
@@ 2.4 Competition log
2.5 Unorganized raw output
A flow of information from the competition not formatted by the platform or online forms needs to be analysed. They can also represent a measure of how active the competition is. The competition forum is the primary source of such information: data exploration posts and notebooks, insights, code sharing, and documentation gleaned on the web on the topic. One can, for example, see an idea appear in a forum post, followed by a code implementation (not necessarily by the same person), followed by a general increase in the scores on the leaderboard. However, relevant technical discussions may happen outside the competition forum, such as on social networks, blog posts, or arXiv papers. Also, competitions are often used as practical projects for (e.g., data science) courses. Compared to typical participants, students are usually compelled to write a hopefully clear document about their techniques, which might be public. Such spontaneous output should be harvested regularly, which is much easier if the competition has a unique acronym to be googled and for which alerts can be set up.
3 Post-challenge workshops and discussion
One or more post-challenge workshops are often organised as part of a conference or in a dedicated venue [^5] https://autodl.chalearn.org/neurips2019666https://llm-efficiency-challenge.github.io/schedule [^7] https://fair-universe.lbl.gov/. Participating in it is part of the incentive for participants. It allows participants to switch from competition to cooperation. They can discuss their techniques between them and with the application or Machine Learning experts. The connections created during such workshops encourage participants to remain engaged with the competition topic instead of just moving on to a different one. The discussions during the workshops and continued online are crucial to deriving the scientific lessons from the workshops and the avenues for further study and possible future challenges.
@@ 3 Post-challenge workshops and discussion
4 Post challenge paper template
@@ 4.1 Introduction
4.2 Challenge Description
This section should ideally come as a reminder and synthesis of a prior paper on the challenge design. One should first review the problem of the challenge to give a good enough understanding to the targeted audience for the remainder of the paper, particularly highlighting its main dimensions of complexity and variations. One should further present the specific task formulated for the challenge with some description of the underlying datasets, framework (such as a chatbot or reinforcement learning framework if relevant) and problem modelisation (possibly relying on some simulator). One can remind if the challenge aims to deliver a new problem or formulation to the Machine Learning community or advance and benchmark state-of-the-art. A section on related works and similar challenges can be written to best position this challenge in the scientific landscape and possibly build upon previous work.
The scoring metric and protocol should be discussed, and some considerations should be shared as to why they were chosen a priori to evaluate any advances towards problem-solving best. A brief description of the challenge platform , as well as possible specific choices (like resource allocation) or developments that can be justified there.
@@ 4.3 Challenge Narrative
For example, Fig. 1 shows the progress of participants in one competition as a function of time; one can see the bulk of participants progressing as a swarm, following community understanding of the problems, while a few isolated outliers obtained the best scores.
Figure 1: Participants’ best score evolution as a function of days in the competition (Amrouche et al., 2019).
@@
Fig. 2 shows participants’ progress in a different competition in the (accuracy, speed) plan. Various strategies are evident: some tried to optimise both simultaneously, why some others (the best), fastrack and gorbuno have first reached the best accuracy, then improved their speed without accuracy loss.
@@
Figure 2: Participants score evolution: the horizontal axis is the accuracy, and the vertical axis is the inference speed. The total score, a function of both variables, is displayed in grey contours. Each colour/marker type corresponds to a contributor; the lines help to follow the score evolution(Amrouche et al., 2021).
@@
A graph like Fig. 3 can help summarise the main events and competition activity.
@@
Figure 3: A high-level retrospective competition period timeline can help understand the competition organisation through phases and noteworthy events such as competition adjustments, peak of activities, performance outbreaks, and collaborative periods. This can support the competition narrative. (Marot et al., 2021)
4.4 Post challenge checks
@@
In most cases, the performance of submissions can be evaluated against a held-out (or private) dataset, providing the final ranking. Generally, this is automatically done on the challenge platform, but it may require some manual actions, mainly when doing it automatically requires too many resources. The stability of the ranking can then be evaluated as in Fig. 4. The graph shows that the rankings among teams are stable, except for 4th Rek, which performs significantly worse in the second phase.
@@
Figure 4: Overfitting/Ranking stability plot from (Xu et al., 2021). The comparison of the x/y axes shows overfitting. In a challenge of two phases, the feedback (or public) phase is the first phase, and the private phase is the second. By showing how submissions perform in the consequent two phases, we demonstrate the overfitting of algorithms. The rectangle shows ranking stability around each team. This rectangle is calculated by the average ranks of multiple reproductions of submissions’ performances.
@@
A more detailed analysis is also possible, as in Fig. 5. The comparison of the private curves shows that a numerical analysis confirms that Gabor (the winner) is clearly above the others. The comparison of private and public curves of different participants shows a clear overfitting case for Lubozs, who was first on the final public leaderboard but slipped to seventh position in the private leaderboard. It turned out that he had indicated in his blog that he had set up an automatic cron job, which was automatically re-submitting new submissions (five times per day, which was the maximum allowed by the platform) with slightly altered parameters to maximise his public score. This engineering feat allowed him to select a lucky spike and grab the top of the public leaderboard, but this did not fool the private leaderboard measurement. It should be noted that this piece of information was not reported through the standard channels or forum; it was discovered by googling, illustrating the point of analysing unorganised information as discussed in Sec. 2.5.
@@
Figure 5: AMS Significance, a figure of merit of a classifier in a physics particle discovery context, as a function of decision threshold for different submissions in the HiggsML challenge (Adam-Bourdarios et al., 2014). The submission score is the maximum of the curve. (d) shows the private curves for different participants. (a),(b) and (c) compare the public and private curves from three participants, Gabor, Choko and Lubozs, with dots indicating their maximum values. The private curves are smoother because they are evaluated using more examples.
@@
Additional graphs can be produced from by-products of solution submission gathered by the platform. For example, Fig. 6 compares the performance on different datasets, giving for each the intrinsic and modelling difficulties. Ideally, organisers should choose datasets of low intrinsic difficulty and high modelling difficulty, which would be in this example Isaac2, Caucase or freddy.
@@
Figure 6: Measurement of task difficulty from (Liu et al., 2020). Each row corresponds to one dataset. Two difficulties are calculated here: intrinsic difficulty (maximum score minus best participant’s score), shown with the green bar and modelling difficulty (best participant’s score minus baseline score), shown with the orange bar.
@@
One can also analyse the solutions to evaluate their originality. For example, Fig 7 shows the dendrogram of a clustering competition. The dendrogram shows how similar or not the clusters found are, which correlates well with what is known on the participant’s method or background: the set of six participants at the top of the diagram (#12#12\#12# 12 to #9#9\#9# 9) have highly optimised the starting kit using generic clustering algorithms; #3#3\#3# 3 and #4#4\#4# 4 are domain experts with the same background, who have found similar clusters with optimised domain methods; #1#1\#1# 1 is a CS student who has spent a lot of time studying the domain literature, has seen similar clusters as #3#3\#3# 3 and #4#4\#4# 4; on the other hand, the diagram sets apart #2#2\#2# 2 who is a CS expert who has developed a very original approach (which turned out to be impractical because of the significant resource it requires).
Figure 7: Dendrogram of the best solutions submitted by participants (name and final rank indicated) to the TrackML challenge (courtesy of authors of (Amrouche et al., 2019)). The diagram shows the thirteen best participants, plus participants ranked 20th, 50th, 100th, as well as the starting kit.
4.5 Deeper analysis of the submissions